Mole of Ar atoms moving at 650 m/s


1
Expert's answer
2022-04-07T22:41:04-0400

Molar mass of Argon, Ar = 39.95 g

Mass of 1 atom, m = 39.95 / (6.023 × 1023) = 6.64 × 10-23 g = 6.64 × 10-26 kg

Now, Kinetic Energy, KE = (1/2) × m × v2 where velocity, v = 650 m/s and mass of 1 atom, m = 6.64 × 10-26 kg

KE = (1/2) × 6.64 × 10-26 kg × (650 m/s)2 = 1.4 × 10-20 J